Solve the following equation:
x/5-x/3=3/2

Answer by stanbon(75887)

x/5-x/3=3/2
---
Multiply thru by 30 to get:
6x - 10x = 45
-4x = 45
x = -45/4
================
Cheers,
Stan H.